OpenStudy (anonymous):

What radius of a circle would guarantee that the circumference is equal to the perimeter of a regular octagon with sides of 6? Use 3.14 for Pi.

OpenStudy (mertsj):

What is the perimeter of the octagon?

OpenStudy (anonymous):

(r=radius of circle) 2*r*pie= 6*8 (perimeter of octagon). then solve from there.

OpenStudy (anonymous):

It doesn't say...But the answer choices are: A)5.5 B)6.4 C)7.6 or D)8.2

OpenStudy (waheguru):

the total premeter of the octagon is 48

OpenStudy (anonymous):

2*r*pie= 6*8 then r=24/pie which equals C) 7.6.

OpenStudy (mertsj):

Oh for heaven's sake. Why don't you guys let the asker answer the questions of someone who is trying to help her understand?

OpenStudy (waheguru):

Yea someone please explain evenn I dint egt this one

OpenStudy (anonymous):

Thank you! :) but I still don't get how you got it lol

OpenStudy (waheguru):

I also got c) 7.6 and can explain if you need help @purenova

OpenStudy (waheguru):

We can use a formula to find this answer First of all we know that the total premeter of the octagon is 6*8=48 Now we use the circumfrence to find the answer the total circumfrence has to be 48 so 2pi r=48 --- solve for r 6.28 * r = 48 48/6.28=r YAAAY